a.)

 N2H4 and H2O2 react according to the equation below.

 N2H4(g)+2H2O2(g)--> N2(g)+ 4H2O(g)

Using the information given in the table, determine the accepted enthalpy of reaction for the following reaction:

Delta H*/KJmol-1  N2H4(g)= +75 H2O2(g)=-133 H2O(g)=-242



b.)

some mean bond enthalpies are given below, based on the enthalpy of reaction calculated in part a, what is the mean bond enthalpy in kj/mol for the O-O bond?

Mean bond enthalpy in KJ/mol-1 N-H=388 N-N=163 N=N=944 H-O=463


2 H--O--O--H --> N=N+4 H--O--H


1
Expert's answer
2021-02-10T01:46:50-0500

a) ∆H = Hf - Hi.

. = -244 - (+75- 133)

= -186KJmol-1

b) Mean bond enthalpy = (388 +163 + 944 +463 )KJmol-1./ 4

= 489.5KJmol-1